OpenStudy (anonymous):

how do i solve a question like this because i didnt get it right on my study guide?? @Shay17

OpenStudy (anonymous):

What's the question? Maybe I can help :)

OpenStudy (anonymous):

What is it @keke_luvs_u

OpenStudy (anonymous):

whoops You mix the letters M,A,T,H,E,M,A,T,I,C,A AND L thoroughly.Without looking, You draw one letter.Find the probability P(A). Write the probability as: a.a fraction in simplest form b.a decimal c.a percent

OpenStudy (anonymous):

How many letters are there in total ? How many A's are there in total ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Ok 4 A's and how many letters in all ?

OpenStudy (anonymous):

12 + the L so 13

OpenStudy (texaschic101):

number of A's = 3 total letters = 12 3/12 = 1/4 = 0.25 = 25%

OpenStudy (anonymous):

@texaschic101 i think im understandning

OpenStudy (anonymous):

That would be true if there were 3 A's and 12 letters but i counted 4 A's and 15 letters

OpenStudy (anonymous):

@keke_luvs_u are you counting the letters in the word AND as well ?

OpenStudy (texaschic101):

M,A,T.H,E,M,A,T,I,C,A,L --- 3 A's and 12 letters

OpenStudy (texaschic101):

I don't think so shay...I didn't count AND...not even sure why it is capitalized

OpenStudy (anonymous):

i ddint either

OpenStudy (anonymous):

Yea that's why I was confused it shouldn't be capitalized if you're not counting it

OpenStudy (texaschic101):

true...capitalizing it makes it confusing

OpenStudy (anonymous):

Well based off the word MATHEMATICAL alone then @texaschic101 is 100% correct !

OpenStudy (texaschic101):

confirmation is always good :)

OpenStudy (anonymous):

Good luck

OpenStudy (anonymous):

sorry about the captilized word

OpenStudy (anonymous):

wait guys can we do 3 more?? that i missed but dont understand?

OpenStudy (anonymous):

Sure post them as new questions though